OpenStudy (anonymous):

use complete sentense to explain the process of graphing f(x)=-x^3+1

OpenStudy (anonymous):

@whpalmer4 @whydoihavetosignupl @EffImHighOffSkittles @eaileen @R00nnYBraiNsbiG @terenzreignz @uri @IloveYou* @primeralph @prettyboy20202 @ajprincess @AnElephant @SnuggieLad @snorl4x @Directrix @dan815 @divagirl421 @DarkKat @Dmitt @faix95 @GOODMAN @hobbs978 @jhonyy9 @Jack17 @KawasumiKimiko @Luigi0210 @Loujoelou @love_jessika15 @lkkk @Callisto @countonme123 @ChrisJP @nincompoop @mukushla @marylou004 @mlb004 @MissHN182 @thefon19 @ryanjnfish @ladycomical @DDCamp @ryanjnfish @jd1028 @phantomofthemathwork @jtrich13 @ajprincess @PaulaCamacho @laxman03 @beccabee77 @babyslapmafro @whpalmer4 @jtrich13 @resident2013 @sumi29 @R00nnYBraiNsbiG @tfuller @Frostbite @SnuggieLad @raque766 @physicsboffin @Falco276 @e.cociuba @blurbendy @kausarsalley @sakigirl @RANE @Danny_Boy @SomeGirl1999 @Jamierox4ev3r @iforgot @lilkittykat36 @n14r96 @ChanceBBrick @knowel @meldo16 @Helpmeplease99 @Manda_luvsu @baseballer20142 @jhalt @RIUKUS @harpelletjain @Math81 @Frank @WildKat @maurimath @shamim @AliaKnight @MayMay_69 @amyamanda @victoria650 @Social @Whatever1 @xxAshxx @lurrch @djoutten @musicalcat98 @Obviously717 @ihategeometry2425 @Zoe_Bailey @RIUKUS @henendez @shaneetta @niah2411 @harpelletjain @Math81 @WildKat @eidan @Hassan07 @Asuman @Frank @bani2100 @turtlebuddy @maurimath @lizzy89 @dborder @peush @EricaaaaB @yehonatan @tori45 @Ojasmin @Markbruce @april115 @RH @shamim @Ryaan @emilyhaddad @Jamierox4ev3r @RANE @NoelGreco @shkrina @lakenclark @Fazeelayaz @sarahusher @JazzyFrizo16 @music101 @HaperFink22 @whydoihavetosignup1 @domu @ashleyjade @Gikundab @Will.Neko @QuinterrianW_23 @tillman--- @baseballer20142 @Manda_luvsu @Charollete @amyamanda @harpreetsk @Whatever1 @HaydenBurden1996 @xxAshxx @lurrch @djoutten @musicalcat98 @Obviously717 @ihategeometry2425 @shaneetta @DLS @NoelGreco @emilyhaddad @marylou004 @JazzyFrizo16 @music101 @HaperFink22 @Courage. @KingAditya @bradely @Gikundab @Will.Neko @QuinterrianW_23 @mcordray @_anastasia0 @ozhobbits @gabigurl @musicalcat98 @Obviously717 @annemoreno @glamourariss @jbsharpe @WildKat @turtlebuddy @turtlebuddy @peush @yehonatan @tori45 @debrajane @EricaaaaB @linabeanax3 @Ojasmin @Markbruce @confusao @sircrashalot @iizabell @tori45 @yehonatan @peush @maurimath @jbsharpe @glamourariss @annemoreno @Obviously717 @mcordray @_anastasia0 @AppylPye @gabigurl @Imtiaz7 @fringe @khadeeja @jhalt @itsmeeQUENNIE @KingAditya @sambhav__jain @whydoihavetosignup1 @mlb004 @Yasminelove1 @caozeyuan @marylou004 @genius12 @agent0smith @AnElephant @lakenclark @JazzyFrizo16 @HaperFink22 @music101 @danielle02 @HaydenBurden1996 @xxAshxx @Obviously717 @Zoe-bailey @peush @yehonatan @tori45 @debrajane @EricaaaaB @linabeanax3 @Fabien1 @Ojasmin @sircrashalot @ankit042 @caozeyuan @rperez36 @freddy_eighty7 @Charollete @freddy_eighty7 @toolCoolChris @agent0smith @genius12 @uri @walac @Luigi0210 @kelliegirl33 @nincompoop @mathmate @stgreen @henpen @binary3i @ChrisJP @zpupster @ankit042 @caozeyuan @betterwiththelightsoff99 @Yasminelove1 @HaperFink22 @glitterythings @whydoihavetosignup1 @Charollete @bradely @ashleyjade @Alexis2332 @Will.Neko @walac @Gikundab @Imtiaz7 @Dianatorres @Alyssa_May_De_Jesus @baseballer20142 @victoria650 @lakenclark @zpupster @mathmate @binary3i @austinL @Loser66 @GoldPhenoix @e.cociuba @DLS @razor99 @Mashy @Machida @RANE @Goten77 @lncognlto @Fifciol @ivancsc1996 @kenbaby @kryton1212 @ankit042 @Henry.Lister @JazzyFrizo16 @betterwiththelightsoff99 @mlb004 @ivano_logy @xxkerstie.bbyxx @MayMay_69 @KawasumiKimiko @itsmeeQUENNIE @Shadly @Albany_Goon @10thgrader @xxmarcusxx1994 @jessy1289 @baseballer20142 @groges @pdd21 @claringleek @Rowa @daniellecrook @imcool @harpreetsk @angelj @JJB @OpenSessame @xseahaven @Whatever1 @harliedanielfrye @borimexz @Jdubose2 @BrianaShamwell @freddy_eighty7 @Horush @snewton29 @Fabien1 @Peter_szlapak @rmpjingwei @mtalampas21 @andriani @mygumshurt @lamel1993 @LittleAnnn @psychomath @yehonatan @heights @JupitersBrain @slayer00647 @2old4math @munastar @jmensal @TanteAnna @Matt-Teepell @DJova @quietkat @bcrisco @larlar77 @Markevia @utvet @felixt03 @9995950976 @natyxz @lftyoutofthepen @kiah44 @Henry.Lister @helder_edwin @Loser66 @inkyvoyd @mathmate @hartnn

OpenStudy (falco276):

Tagging to many people on OS is considered spamming. Otherwise if you keep on doing this, it will suspend your account. I have been warned once not to do this. Just letting you know. :)

OpenStudy (anonymous):

i didnt do that i have no idea who did

OpenStudy (anonymous):

i was thinking the same thing to many people

OpenStudy (anonymous):

ive been tagged like more than 50 times

OpenStudy (anonymous):

Open your heart to Jesus Christ.

OpenStudy (anonymous):

i just need help on this one question can anyone help

OpenStudy (falco276):

I'll do. What is it? :)

OpenStudy (anonymous):

thank you ut is: The binomial x + 4 is the divisor of 3x3 + 5x2 – 26x + 8. Use complete sentences to explain how you use (either) polynomial division or synthetic division to find the quotient. Be sure to provide this quotient in your answer.

OpenStudy (abb0t):

We know you did it @barrelracer1234 and now i will suspend you!

OpenStudy (falco276):

Algebra? Am horrible. Sorry. :(

OpenStudy (falco276):

@abb0t Don't suspend him. He's innocent.

OpenStudy (anonymous):

it is ok look at who posted t @abb0t it said someone by the user xfdjerdjyfvd

OpenStudy (anonymous):

and i dont apperciate being accused of something i did not do all i did was ask one question that i don not understand

OpenStudy (falco276):

Barrelracer1234 is right. I agree with him.

OpenStudy (anonymous):

im a girl btw

OpenStudy (falco276):

never mind.

OpenStudy (anonymous):

i am not the one who posted all those names if you look you can see it was not me who posted it

OpenStudy (anonymous):

ok i see that but i am not the one who posted it so i am innocent

OpenStudy (anonymous):

all i did was ask for help on a question i do not understand and all you are doing is accusing me of something i did not do and i dont apperciate that

OpenStudy (anonymous):

i will really would like some help on this one question so i dont have to bother anyone else for the day

OpenStudy (anonymous):

can you please help me @claringleek

OpenStudy (anonymous):

Yea sure :)

OpenStudy (anonymous):

Let x^(1/3) = u ? x^(2/3) = u? Substitute into equation and solve for u ? 3u? + u - 2 = 0 ? 3u? + u - 2 = 0 ? 3u? + 3u - 2u - 2 = 0 ? 3u (u+1) - 2 (u+1) = 0 ? (3u - 2) (u+1) = 0 ? (3u - 2) = 0 or (u + 1) = 0 ? u = 2/3 or -1 From original [ [ substitution x^(1/3) = u ? x^(1/3) = u ? x = u? ? x = (2/3)? or (-1)? ? x = 8/27 or -1 ] ]

OpenStudy (anonymous):

Thats what i know but im not suree how to put it in a sentence im still trying though

OpenStudy (anonymous):

ok thanks so much i understand this now

OpenStudy (anonymous):

lol ok :) its alot yes but i gets easy lol i feel like such a nerd :)